# Jon Danielsson London School of Economics Andreas Uthemann

### 10 Jan 2024 v3

@@ 2.3 System responses
A key challenge for AI working for the macroprudential, but not generally for the microprudential authorities, relates to the dynamic interaction of the financial system participants. A helpful framework for understanding the problem is the Lucas (1976) critique, which states that the decision rules used by economic agents, financial institutions in our case, depend on the underlying economic environment, and can change as regulations change undermining their effectiveness. An example of Goodhart’s (1974) law, “Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes.” Changes to financial regulations or the level of supervision, including changes to the crisis resolution playbook, as well as decisions taken during resolution, in response to a given crisis will change the responses of the private sector to a similar crisis in the future in unexpected ways.
Of particular interest when AI is to be used for economic analysis relates to how Lucas and Goodhart came to their conclusions. Prior to the 1970s, economists believed in large general models informed by statistical patterns, with the Phillips curve, representing a trade-off between unemployment and inflation based on a negative empirical correlation of these two variables, as a centrepiece. [^2] For an important example, see Rasche and Shapiro (1968) explaining the MIT-Fed model used by the Federal Reserve and developed in collaboration with MIT. The contribution of Lucas was to recognise that when the authorities tried to exploit this apparent trade-off, they changed the way the private sector negotiated wages and set prices, leading to a breakdown of the previously identify correlations in the data. Further attempts to stimulate the economy through this channel created inflation without reducing unemployment giving rise to the word stagflation. In response to this lesson, the economic profession moved away from large-scale simultaneous equation models based on statistical correlations to smaller, structural models that explicitly modelled beliefs and general equilibrium effects so that reactions to policy interventions followed a well understood and internally consistent logic encoded in the model, perhaps an early example of interpretability.
@@
The relative importance of exogenous versus endogenous risk depends on the problem. For most microprudential regulations, the frequency of events and lack of large strategic interactions means that assuming risk is exogenous is usually quite innocuous. However, any authority using data driven analytics that uses exogenous risk measurement for assessing the risk of financial instability will likely be seriously misled as to how the financial system evolves in times of stress.
@@ 2.4 Pre-specified objectives and distributed decision making
Mutable objectives do not pose much of a problem to micropru. The rulebook is known, and static on the timescale most decisions are made. It does evolve over time in response to events and the regulated response to regulations. Still, AI can quickly update its understanding of the objectives, both by adopting changes to the rulebook and by learning from observing how the human supervisors act.
@@
This is not the case in macropru, as it operates on very long timescales, the time between relevant events is usually in decades. Furthermore, it can be very difficult to define the macroprudential objective except at the highest levels of abstraction, such as preventing severe dysfunction in key financial markets and especially the failure of systemically important institutions. This applies to the design of macroprudential regulations, macroprudential supervision and crisis resolution. It can be difficult to make a case for the need to allocate significant resources to prevent something that only happens in the distant future, and the macroprudential regulations are susceptible to lobbying and political interference, which means that the objectives of regulations can change over time. Learning from previous crisis interventions might not be of much help as the concrete goals the actions under consideration are to achieve are likely specific to particular circumstances and political environments.
@@
The most severe financial crises can have catastrophic consequences if not addressed adequately, with direct economic costs in the several trillions of dollars, as noted by Barnichon et al. (2022), and a large number of people materially affected. When that happens, society demands we do what it takes to resolve the crisis. Often, the rules and the laws in place might stand in the way of the most effective crisis resolution. Emergency sessions of Parliament to rectify that are not uncommon, such as Switzerland’s resolution of Credit Suisse. [^3] https://www.bloomberg.com/news/articles/2023-03-20/credit-suisse-collapse-reveals-some-ugly-truths-about-switzerland-for-investors. Emergency constution clauses might be invoked. Pistor (2013) in her legal study of the resolution of financial crises, finds that if the existing law prevents the most effective course of action, there is acceptance from the political and judicial system to suspend the law in the name of the higher objective of crisis resolution. Furthermore, when a severe crisis happens, the political leadership takes charge, inevitable because if it becomes necessary to change or bypass the law or significantly redistribute resources, the political leadership is the only entity with the necessary legitimacy. Given the fluidity of this process, it is difficult to see how one would specify the objectives for an AI so that it can provide real-time advice for crisis resolution or make decisions.
@@
We have a long experience of resolving crises and have a relatively good understanding of the process. The regulatory system is usually modular, with separate authorities and fiercely guarded mandates. In the most severe crises, these silos break down. All relevant authorities, the affected private sector, the judiciary and especially the political leadership come together to decide how to resolve the crisis. Government ministers usually lead this process. This may involve the same entities in other financial centres in a global crisis. Each stakeholder brings their own education, ethics, professional experience and objectives to the table. Such a process can be highly robust. All pertinent issues are discussed, including information that was until then confidential or implicit. Such analysis depends on implicit knowledge and intuitive understanding its participants have of the current situation and each others’ views and objectives. Intuition can mean finding analogous problems from seemingly unrelated domains that can provide “create solutions” for the current crisis that had not been encountered in this form in its problem domain. This process of successful extrapolation is often vital to crisis resolution and can imply ignoring or revaluing previously pursued goals. There are many such examples in history. In his re-evaluation of the 1866 crisis, Bagehot (1873) invented modern financial stability policies, like lending of last resort, while the German central banker, Hjalmar Schacht, used a short squeezing to stop the hyperinflation at the time. The Swedish central bank in 1992 created the good bank-bad bank model for crisis resolution. Reproducing this process will be particularly difficult for AI as it implies both shifting attention to seemingly unrelated problems and potentially modifying its objective function.
@@
Crisis resolution is arguably the most important aspect of financial policy, especially for central banks — one of their raisons d’être. Nevertheless, given the above challenges — scarce data, unknown-unknowns, endogenous structural changes in response to attempted control — humans struggle with this task and this is where AI could benefit them the most. Unfortunately, it has to overcome the same conceptual problems and this challenge is as difficult for AI as for human regulators, if not more so in certain dimensions. Paradoxically, progress in AI’s suitability for the task might come from a better human understanding of these problems that can then be translated into better algorithms.
@@ 2.5 Incentives
The objectives of the various market participants are specific to them and often in conflict with those of other participants and society. That is why we have financial regulations: to align the private sector’s incentives with society at large. The interactions between financial supervisors and the private sector decision-makers can be seen as a principal-agent (PA) relationship, where the highly resourced and sophisticated private sector — the agent — wants to get on with successfully running its business while the supervisors — the principal — aims to ensure that its agents operate a financial system that efficiently channels savings to investments while not abusing their clients and not taking excessive risks that can cause costly crises.
@@
The PA problem applied to financial markets is particularly difficult for three reasons. First, contracting is incomplete, meaning we cannot specify what will happen in every contingency, we might not even know what the contingencies are. Second, the very behaviour to be controlled is risk taking, and since risk is a latent variable, it cannot be measured directly but only imprecisely inferred from market movements. Consequently, it can be difficult to determine whether an undesirable outcome is due to misbehaviour or innocent bad luck. Finally, because implicit or explicit government guarantees are fundamental to all financial regulations, market participants, knowing that some of their risk taking is publicly insured, can behave in a way that makes government bailouts more likely, creating moral hazard. [^4] Dewatripont et al. (1994) provide a detailed analysis of the problem of financial regulation with incomplete contracts, government guarantees and private sector moral hazard.
@@
As we increasingly use AI in regulations and the private sector, a new dimension in the PA problem emerges. The one-sided financial institution-regulator problem becomes two-sided, with the addition of the relationship between the regulators and financial institutions as principals and their respective AIs as agents. The reason is that we cannot expect the AI’s objectives to be fully aligned with its human operator as AI can take hidden actions to achieve its goals. This means the human operator needs to incentivise the AI to act in its owner’s interest.
@@
The presence of the two-sided PA problem amplifies the already significant existing PA problem in financial regulation. We have in the past seen many cases where market participants actively and deliberately create heightened stress since anybody forewarned of stress can profit. Hence, creating and amplifying it is profitable. Lowenstein (2000) shows just one example from the LTCM crisis in 1998. AI might find it even easier to act in that way. It will therefore be essential for the authorities to explicitly consider how to align the AI engines, whether directly operated by the regulated or by their AI vendors, with the objectives of the authority.
@@
Existing mechanisms for managing the regulator-bank PA relationship likely will not be adequate when AI is involved. AI commands more sophistication. In the technical terminology of moral hazard, it has more hidden actions at its disposal and will command higher information rents. In addition, AI’s objectives can be less straightforward than those of traditional financial institutions that trade-off risks and returns. How do you incentivise an algorithm whose objective is to correctly predict the next token? What higher level objectives does this loss function induce, and how do we formulate an incentive scheme that imposes the correct constraints on AI behaviour? How can we make it internalise negative externalities such as the possibilities of deliberately creating fire sales or runs to trigger privately beneficial but publicly costly bailouts?
@@ 3 Artificial intelligence and machine learning
There is no standard definition of artificial intelligence (AI). One of the pioneers of AI, John McCarthy, saw it as “the science and engineering of making intelligent machines”, [^5] http://jmc.stanford.edu/artificial-intelligence/what-is-ai/index.html where perhaps “intelligence measures an agent’s ability to achieve goals in a wide range of environments” in the definition of Legg and Hutter (2007). Generally, it is a computer algorithm that performs tasks that would otherwise be done by a human. For a general view of AI, Russel (2019) is particularly useful.
@@
ML provides the fundamental building blocks of most AI, finding statistical regularities in large data sets helping providing AI engines with a deeper understanding of the underlying latent structure that generates the data of a given problem domain. One of the key differences to conventional applied statistics is that the input data tend to be unstructured, are often text-based and large in size, and that the models can encode highly complex statistical relationships. The modeller usually does not supply prior understanding of the parameterised model structure. See the Appendix for details on ML useful in financial applications.
@@ 3.1 From reactive machines to AGI
Even though the ability of AI engines has increased rapidly, today’s AI merely matches patterns and does not understand in the same way a human does. It is unclear what achieving such understanding that will involve. When considering the ability of the various AI, including those only hypothesised, it is beneficial to classify AI based on its abilities. As it turns out, that is surprisingly difficult task since the experts do not agree on classification, terminology, ability or future potential. At the risk of oversimplification, there are four types of AI.
The first is often termed a reactive machine, which is a system with no memory of previous decisions, and designed to perform a specific task. In other words, it it only works with already available data. Typical examples are a classification engines that identify objects and photographs and recommendation engines from vendors.
@@
We then have a theory of mind, not yet realised, AI that aims to understand thoughts and emotions, and consequently simulate human relationships. It can reason and personalise interactions. Importantly from the point of view of financial policy, a theory of mind AI should be able to understand and provide context for policy decisions, which today’s AI cannot do, and hence be invaluable in crises resolution. AI that understands how its interlocutors think can explain its decisions to them, and reasons about how they will react to its actions and plan accordingly.
And finally, artificial general intelligence or AGI. Still purely theoretical while often hypothesised. It would use existing knowledge to learn new things without needing human beings for training, and being able to perform any intellectual task humans can. And if it reaches that level, increased computer capacity would allow it to surpass humans. [^6] Kasirzadeh discusses existential risks arising from AGI
The potential for the eventual emergence of a theory of mind and AGI remains controversial, and experts have strong disagreements on both the likelihood of such eventualities and what it takes to reach them. The ability of an AI engine depends on three factors: network structure, data and compute. The more relevant data we feed into the engine and the more compute we apply, the better it will perform. What is controversial is whether that is all we all we need to make progress on theory of mind and AGI, or if a conceptual leap in the ways of learning is also required.
A point of controversy is whether all that is needed to achieve AGI is next token prediction, in other words, affecting the next element in a sequence, whatever the sequence, in which case all that is required for AGI is sufficient amount of training data and compute, we would not need conceptual improvements in network technology.
When transiting from using AI only to learn about the world to using it to advise on real-world decisions and even make such decisions, then a key challenge is AI alignment and safety, that is, how to make AI systems behave in line with human intentions and values. [^7] For a survey of AI alignment work, see, e.g., Ji et al. (2023), and Shevlane et al. (2023), and Bengio et al. (2023) model evaluation for extreme risks.
@@ 3.2 How do we know AI does what it is supposed to?
A key challenge is that the training algorithms need to know what to optimise for. In other words, they need a well specified objective function that evaluates the cost and benefits of alternative courses of action given the current state of the environment and its future evolution. In most financial applications, they need to take into account how the system reacts to its actions. Misspecifying the problem, or what is likely to be more common, insufficient specification of the problem, leads to suboptimal, and even catastrophically bad, decisions.
In the 1980s, an AI decision support engine called EURISKO used a cute trick to defeat all its human competitors in a naval wargame, sinking its slowest ships to maintain manoeuvrability. This early example of AI reward hacking, something humans are experts in, illustrates how difficult it is to trust AI. How do we know it will do the right thing? Human admirals don’t have to be told not to sink their own ships, and if they do, they either have high-level political acquiescence or are stopped by their junior officers. Any current AI making autonomous decisions has to be told or learn from observing human decisions that sinking its own ships is not allowed. The problem is that the real world is far too complex for us to train AI on every eventuality. AI will predictably run into cases where it will make critical decisions in a way that no human would. EURISKO’s creator, Douglas Lenat, notes that “[w]hat EURISKO found were not fundamental rules for fleet and ship design; rather, it uncovered anomalies, fortuitous interactions among rules, unrealistic loopholes that hadn’t been foreseen” (Lenat, 1983, p 82). Each of EURISKO’s three successive victories resulted in rule changes intended to prevent repetition. Still, in the end, the only thing that worked was telling Lenat that his, and his AI’s, presence was not welcome.
If we ask ChatGPT whether it is okay for admirals to sink their own ships, it says no. However, this is a well known example, likely in the canon ChatGPT trained on. Would it have come up with the same answer if Lenat had not entered EURISKO in the naval board game? We don’t know.
@@
That means it is essential that AI can explain its reasoning, just like humans are required to. That can be difficult since AI’s outputs are simply a nonlinear mapping of input data, where decisions are affected both by the network structure and trained parameters. The machine says yes or no based on one particular configuration of a trillion parameters.
@@
AI that cannot explain its reasoning cannot be used for many applications in finance and elsewhere. Not surprisingly, considerable resources are being brought to bear on that problem, allowing the operators to trace particular recommendations to the input data, so instead of a recommendation being merely a function of model parameters, they are determined by particular inputs that are analysed logically. Benchmarking AI engines to specific tasts could be particularly valuable in fianncial applications as we discuss in Section 6.

### 03 Feb 2024 v4

@@ 2.3 System responses
A key challenge for AI working for the macroprudential, but not generally for the microprudential authorities, relates to the dynamic interaction of the financial system participants. A helpful framework for understanding the problem is the Lucas (1976) critique, which states that the decision rules used by economic agents depend on the underlying economic environment and can change as regulations change, undermining their effectiveness. An example of Goodhart’s (1974) law, “Any observed statistical regularity will tend to collapse once pressure is placed upon it for control purposes.” Changes to financial regulations or the level of supervision, including changes to the crisis resolution playbook, as well as decisions taken during resolution, will change the responses of the private sector to a similar crisis in the future in unexpected ways.
Of particular interest when AI is to be used for economic analysis relates to how Lucas and Goodhart came to their conclusions. Prior to the 1970s, economists believed in large general models informed by statistical patterns, with the Phillips curve, representing a trade-off between unemployment and inflation based on a negative empirical correlation of these two variables, as a centrepiece. [^2] For an important example, see Rasche and Shapiro (1968) explaining the MIT-Fed model used by the Federal Reserve and developed in collaboration with MIT. The contribution of Lucas was to recognise that when the authorities try to exploit this apparent trade-off, they change the way the private sector negotiates wages and sets prices, leading to a breakdown of previously identified correlations. Further attempts to stimulate the economy through this channel create inflation without reducing unemployment, giving rise to the word stagflation. In response, the economic profession moved away from large-scale simultaneous equation models based on statistical correlations to smaller, structural models that explicitly modelled beliefs and general equilibrium effects so that reactions to policy interventions followed a well understood and internally consistent logic encoded in the model, perhaps an early example of interpretability. An AI engine would need to do the same if it were to become an effective macropru regulator.
@@
The relative importance of exogenous versus endogenous risk depends on the problem. For most microprudential regulations, the frequency of events and lack of large strategic interactions means that assuming risk is exogenous is usually quite innocuous. However, any authority using data driven analytics that uses exogenous risk measurement for assessing the risk of financial instability, will likely be seriously misled as to how the financial system evolves in times of stress.
@@ 2.4 Pre-specified objectives and distributed decision making
Mutable objectives do not pose much of a problem to micropru. The rulebook is known and usually static on the timescale decisions are made. While it evolves in response to events and the regulated response to regulations, AI can quickly update its understanding of the objectives, by adopting changes to the rulebook, learning from observing how the human supervisors act and reinforcement learning.
@@
This is not the case in macropru, as it operates on very long timescales. The time between relevant events is usually in decades. It can be very difficult to define the macroprudential objective except at the highest levels of abstraction, such as the prevention of severe dysfunction in key financial markets and especially the failure of systemically important institutions. This applies to the design of macroprudential regulations, macroprudential supervision and crisis resolution. It can be difficult to make a case for the need to allocate significant resources to prevent something that only happens in the distant future, and the macroprudential regulations are susceptible to lobbying and political interference, which means that the objectives of regulations can change over time. Learning from previous crisis interventions might not be of much help as the concrete goals the actions under consideration are to achieve are likely specific to particular circumstances and political environments.
@@
The most severe financial crises can have catastrophic consequences if not addressed adequately, with direct economic costs in the several trillions of dollars, as noted by Barnichon et al. (2022), and a large number of people materially affected. When that happens, society demands we do what it takes to resolve the crisis. Often, the rules and the laws in place might stand in the way of the most effective crisis resolution. Emergency sessions of Parliament to rectify that are not uncommon, such as Switzerland’s resolution of Credit Suisse. [^3] https://www.bloomberg.com/news/articles/2023-03-20/credit-suisse-collapse-reveals-some-ugly-truths-about-switzerland-for-investors. Emergency constitution clauses might be invoked. Pistor (2013), in her legal study of the resolution of financial crises, finds that if the existing law prevents the most effective course of action, there is acceptance from the political and judicial system to suspend the law in the name of the higher objective of crisis resolution. Furthermore, when a severe crisis happens, the political leadership takes charge, inevitable because if it becomes necessary to change or bypass the law or significantly redistribute resources, the political leadership is the only entity with the necessary legitimacy. Given the fluidity of this process, it is difficult to see how one would specify the objectives for an AI so that it can provide real-time advice for crisis resolution or make decisions.
@@
We have a long experience of resolving crises and have a relatively good understanding of the process. The regulatory system is usually modular, with separate authorities and fiercely guarded mandates. In the most severe crises, these silos break down. All relevant authorities, the affected private sector, the judiciary, and especially the political leadership have come together to decide how to resolve the crisis. Government ministers usually lead this process. This may involve the same entities in other financial centres in a global crisis. Each stakeholder brings their own education, ethics, professional experience and objectives to the table. Such a process can be highly robust. All pertinent issues are discussed, including information that was confidential or implicit until then. Such analysis depends on implicit knowledge and intuitive understanding its participants have of the current situation and each others’ views and objectives. Intuition can mean finding analogous problems from seemingly unrelated domains that can provide creative solutions for the current crisis that has not been encountered before. This process of successful extrapolation is often vital to crisis resolution and can imply ignoring or revaluing previously pursued goals. There are many such examples in history. In his re-evaluation of the 1866 crisis, Bagehot (1873) invented modern financial stability policies, like lending of last resort, while the German central banker, Hjalmar Schacht, used a short squeezing to stop the hyperinflation in 1923. The Swedish central bank in 1992 created the good bank-bad bank model for crisis resolution. Reproducing this process will be particularly difficult for AI as it implies both shifting attention to seemingly unrelated problems and potentially modifying its objective function.
@@
Crisis resolution is arguably the most important aspect of financial policy, especially for central banks — one of their raisons d’être. Nevertheless, given the above challenges — scarce data, unknown-unknowns, endogenous structural changes in response to attempted control — humans struggle with this task, and this is where AI could benefit them the most. Unfortunately, it has to overcome the same conceptual problems, and this challenge is as difficult for AI as for human regulators, if not more so. Paradoxically, progress in AI’s suitability for the task might come from a better human understanding of these problems that can then be translated into better algorithms.
@@ 2.5 Incentives
The objectives of the various market participants are specific to them and often in conflict with those of other participants and society. That is why we have financial regulations: to align the private sector’s incentives with society at large. The interactions between financial supervisors and the private sector decision-makers can be seen as a principal-agent (PA) relationship, where the highly resourced and sophisticated private sector — the agent — wants to get on with successfully running its business. In contrast, the supervisors — the principal — aim to ensure that the agents operate a financial system that efficiently channels savings to investments while not abusing their clients or taking excessive risks that can cause costly crises.
@@
The PA problem applied to financial markets is particularly difficult for three reasons. First, contracting is incomplete, meaning we cannot specify what will happen in every contingency. We might not even know what the contingencies are. Second, the very behaviour to be controlled is risk taking, and since risk is a latent variable, it cannot be measured directly but only imprecisely inferred from market movements. Consequently, it can be difficult to determine whether an undesirable outcome is due to misbehaviour or innocent bad luck. Finally, because implicit or explicit government guarantees are fundamental to all financial regulations, market participants, knowing that some of their risk taking is publicly insured, can behave in a way that makes government bailouts more likely, creating moral hazard. [^4] Dewatripont et al. (1994) provides a detailed analysis of the problem of financial regulation with incomplete contracts, government guarantees and private sector moral hazard.
@@
As we increasingly use AI in regulations and the private sector, a new dimension in the PA problem emerges. The one-sided financial institution-regulator problem becomes two-sided, with the addition of the relationship between the regulators and financial institutions as principals and their respective AI as agents. The reason is that we cannot expect the AI engines’ objectives to be fully aligned with its human operator as AI can take hidden actions to achieve its goals. This means the human operator needs to incentivise the AI to act in its owner’s interest.
@@
The presence of the two-sided PA problem amplifies the already significant existing PA problem in financial regulation. We have in the past seen many cases where market participants actively and deliberately create heightened stress since anybody forewarned of stress can profit. Creating and amplifying crises is profitable. Lowenstein (2000) shows just one example from the LTCM crisis in 1998. AI might find it even easier to act in that way. It will, therefore, be essential for the authorities to explicitly consider how to align the AI engines, whether directly operated by the regulated or by their AI vendors, with the objectives of the authority.
@@
Existing mechanisms for managing the regulator-bank PA relationship will not be adequate when AI is involved. In the technical terminology of moral hazard, AI has more hidden actions at its disposal and will command higher information rents. In addition, AI’s objectives can be less straightforward than those of traditional financial institutions that trade-off risks and returns. How do you incentivise an algorithm whose objective is to predict the next token correctly? What higher level objectives does this loss function induce, and how do we formulate an incentive scheme that imposes the correct constraints on AI behaviour? How can we make it internalise negative externalities, such as the possibility of deliberately creating fire sales or runs to trigger privately beneficial but publicly costly bailouts?
@@ 3 Artificial intelligence and machine learning
There is no standard definition of artificial intelligence (AI). One of the pioneers of AI, John McCarthy, saw it as “the science and engineering of making intelligent machines”, [^5] http://jmc.stanford.edu/artificial-intelligence/what-is-ai/index.html where perhaps “intelligence measures an agent’s ability to achieve goals in a wide range of environments” in the definition of Legg and Hutter (2007). Generally, it is a computer algorithm that performs tasks that a human would otherwise do. For a general view of AI, Russel (2019) is particularly useful.
@@
ML provides the fundamental building blocks of most AI, finding statistical regularities in large data sets and helping provide AI engines with a deeper understanding of the underlying latent structure that generates the data of a given problem domain. One of the key differences to conventional applied statistics is that the input data tend to be unstructured, are often text-based and large in size, where the models can encode highly complex statistical relationships. The modeller usually does not supply a prior understanding of the parameterised model structure. See the Appendix for details on ML that are useful in financial applications.
@@ 3.1 From reactive machines to AGI
Even though the ability of AI engines has increased rapidly, today’s AI merely match patterns and do not understand in the same way a human does. It is unclear what achieving such an understanding will involve. When considering the abilities of various AIs, including those only hypothesised, it is beneficial to classify AI based on its abilities. As it turns out, that is a surprisingly difficult task since the experts do not agree on classification, terminology, ability or future potential. At the risk of oversimplification, there are four types of AI.
The first is often termed a reactive machine, which is a system with no memory of previous decisions and designed to perform a specific task. In other words, it only works with already available data. Typical examples are classification engines that identify objects and photographs and recommendation engines.
@@
We then have a theory of mind, not yet realised, AI that aims to understand thoughts and emotions and consequently simulate human relationships. They can reason and personalise interactions. Importantly, from the point of view of financial policy, a theory of mind AI should be able to understand and provide context for policy decisions, which today’s AI cannot do, and hence be invaluable in crisis resolution. AI that understands how its interlocutors think can explain its decisions to them and reason about how they will react to its actions and plan accordingly.
And finally, artificial general intelligence or AGI. It is still purely theoretical while often hypothesised. It would use existing knowledge to learn new things without needing human beings for training and being able to perform any intellectual task humans can. And if it reaches that level, increased computer capacity would allow it to surpass humans. [^6] Kasirzadeh discusses existential risks arising from AGI
The potential for the eventual emergence of a theory of mind and AGI AI remains controversial, and experts have strong disagreements on both the likelihood of such eventualities and what it takes to reach them. The ability of an AI engine depends on three factors: network structure, data and compute. The more relevant data we feed into the engine and the more compute we apply, the better it will perform. What is controversial is whether that is all we need to make progress on the theory of mind and AGI or if a conceptual leap in the ways of learning is also required.
A point of controversy is whether the next token prediction is sufficient for achieving AGI. In that case, all that is required for AGI is a sufficient amount of training data and compute. We would not need conceptual improvements in network technology.
When transiting from using AI only to learn about the world to use it to advise on real-world decisions and even make such decisions, a key challenge is AI alignment and safety, that is, how to make AI systems behave in line with human intentions and values. [^7] For a survey of AI alignment work, see, e.g., Ji et al. (2023), and Shevlane et al. (2023), and Bengio et al. (2023) model evaluation for extreme risks.
@@ 3.2 How do we know AI does what it is supposed to?
A key challenge is that the training algorithms need to know what to optimise for. In other words, they need a well specified objective function that evaluates the cost and benefits of alternative courses of action given the current state of the environment and its future evolution. In most financial applications, they need to take into account how the system reacts to its actions. Misspecifying the problem, or what is likely to be more common, insufficient specification of the problem leads to suboptimal and even catastrophically bad decisions.
In the 1980s, an AI decision support engine called EURISKO used a cute trick to defeat all its human competitors in a naval wargame, sinking its slowest ships to maintain manoeuvrability. This early example of AI reward hacking, something humans are experts in, illustrates how difficult it is to trust AI. How do we know it will do the right thing? Human admirals don’t have to be told not to sink their own ships, and if they do, they either have high-level political acquiescence or are stopped by their junior officers. Any current AI making autonomous decisions has to be told, or learn from observing human decisions, that sinking its own ships is not allowed. The problem is that the real world is far too complex for us to train AI on every eventuality. AI will predictably run into cases where it will make critical decisions in a way that no human would. EURISKO’s creator, Douglas Lenat, notes that “[w]hat EURISKO found were not fundamental rules for fleet and ship design; rather, it uncovered anomalies, fortuitous interactions among rules, unrealistic loopholes that hadn’t been foreseen” (Lenat, 1983, p 82). Each of EURISKO’s three successive victories resulted in rule changes intended to prevent repetition. Still, in the end, the only thing that worked was telling Lenat that his and his AIs’ presence was not welcome.
If we ask ChatGPT whether it is okay for admirals to sink their own ships, it gives a well-grounded, nuanced answer. However, this is a well known example, likely in the canon ChatGPT trained on. Would it have come up with the same answer if Lenat had not entered EURISKO in the naval board game? We don’t know.
@@
That means it is essential that AI can explain its reasoning, just like humans are required to. That can be difficult since AI outputs are simply a nonlinear mapping of input data, where decisions are affected both by the network structure and trained parameters. The machine may say yes or no based on one particular configuration of a trillion parameters.
@@
AI, which cannot explain its reasoning, cannot be used for many applications. Not surprisingly, considerable resources are being brought to bear on that problem, allowing the operators to trace particular recommendations to the input data, so instead of a recommendation being merely a function of model parameters, they are determined by particular inputs that are analysed logically. Benchmarking AI engines to specific tasks could be particularly valuable in financial applications, as we discuss in Section 6.
4 AI use in the financial system
Many public and private financial system applications can be supported by the two forms of AI accessible today, reactive and limited memory. And, as the engines’ capabilities develop at a rapid pace, and we come closer to a theory of mind, and possibly AGI, we expect AI use to expand in both the public and private sectors (Maple et al., 2023; Kinywamaghana and Steffen, 2021; Korinek, 2023).
Credit allocation, fraud, compliance and risk management are just a few of the tasks in which reactive machines have proven highly valuable. Supervisors may use them to develop the rulebook and enforce compliance with it. There is no technical reason why AI cannot perform the majority of such jobs now; what is holding it back is a lack of human expertise, compute and data sharing (see Section 2.1) as well as caution.
For private financial institutions, where senior managers can be held to account, the unknown territory of AI, compared to relatively safe and familiar incumbent technology, might represent unacceptable personal risk. However, while legacy systems might hold back traditional financial institutions and perhaps a culture that reflects traditional practices, challenger banks find it much easier to use technology. In a competitive market, their greater efficiency and better delivery of financial services will most likely push AI adoption across the private sector.
The authorities have chosen a more cautious approach to AI. However, as the private sector develops its use of AI, the authorities must follow suit if they are to remain relevant. They are already doing so, as evidenced by Moufakkir (2023) and Cook (2023). Anecdotal evidence suggests that current AI use is mostly limited to regular forecasting and low-level data processing. However, most authorities are only now beginning a more structured analysis of AI.
4.1 Authorities taking advantage of AI
While financial authorities have so far used AI fairly sparingly, we expect that to change. This will necessitate investments in data, human capital and computing power. The most straightforward is data, as authorities already have access to large databases, including confidential disclosures, that can be used to train AI. This includes:
  • Observations on past compliance and supervisory decisions
  • Prices, trading volumes, and securities holdings in fixed-income, repo, derivatives and equity markets
  • Assets and liabilities of commercial banks
  • Network connections, like cross-institution exposures, including cross-border
  • Textual data
    • The rulebook
    • Central bank speeches, policy decisions, staff analysis
    • Records of past crisis resolution
  • Internal economic models
    • Interest rate term structure models
    • Market liquidity models
    • Inflation, GDP and labour market forecasting models
    • Equilibrium macro model for policy analysis
The authorities will find human resources and compute more difficult, not the least because of very high costs of human capital specialised in AI, as we discuss below in Section 5.4. Training AI engines is very expensive and necessitates access to specialised facilities, both of which are out of reach for most financial authorities. However, because many of the largest costs are associated with training general-purpose engines, it is easy to overstate those problems. The marginal cost to the authorities is likely to be significantly smaller, as they may be able to extend extant engines via transfer learning on curated datasets for financial sector use that include the items listed above.
There are many activities within the financial authorities where AI would be of considerable help. Start with the microprudential regulations, both design and supervision. ML can be used to translate the rulebook into a logic engine, allowing analysis of internal consistency and coverage and facilitating AI supervision. Several authorities have already made progress in this area. Generative models can simulate undesirable behaviour and hence help in designing the rulebook to protect against such outcomes. Reactive engines can identify undesirable behaviour not covered by extant rules and propose remedies.
AI can be beneficial in supervision, providing routine compliance advice to the private sector, recommending regulatory actions, and even making supervisory decisions. While human supervisors would initially closely oversee it, reinforcement learning with human feedback will help supervisory AI to become increasingly performant. For challenging areas of microprudential supervision, adversarial architectures such as generative adversarial networks (GANs) might be particularly beneficial in understanding complex areas of authority-private sector interactions, such as fraud detection.
It will also likely be useful in other activities, such as ordinary economic analysis and forecasting, thanks to an economic engine that performs both theoretical and statistical analysis. This could be achieved with a general purpose foundation model augmented via transfer learning using publicly and privately available data, established economic theory embodied in economic models and previous policy analysis. Reinforcement learning with feedback from human experts might be useful to improve the engine further. Such AI would be very beneficial to those conducting macroprudential stress tests. A key challenge in such analysis (Anderson et al., 2018) is that the stress needs to cover behaviour that we don’t usually see in the system and capture particular interrelations between the various market participants in times of stress. Here, generative models could be valuable in facilitating the running of scenarios in crisis resolution, helping to identify and analyse the drivers of extreme market stress. The authorities could also use such models as artificial labs to experiment on policies and evaluate private sector algorithms, perhaps building on existing agent-based simulations. Over time, as AI technology improves, theory of mind AI might be able to understand and provide context for policy decisions and hence be invaluable in crisis resolution. AI that understands how its interlocutors think can explain its decisions to them, as well as reason about how they will react to its actions and plan accordingly, will be of considerable benefit. Not the least, it could address the problem of misspecified objectives.
4.2 The risk of authorities losing control
4.3 Trust
While an authority might not wish to get to that point, its use of AI might end up there regardless. As we come to trust AI analysis and decisions and see how well it performs in increasingly complex and essential tasks, it may end up by stealth, where the authorities do not want to be. That arises because of how AI creates trust. There are many examples where technology was initially met with scepticism. As it performs, technology is then increasingly accepted, as we have seen with AI landing passenger aeroplanes, AI-guided surgery and self-driving cars. Even if we only allow AI to execute the simplest tasks inside the financial authorities, as it starts performing at the same level or better than its human counterparts, but at a much lower cost, we can expect AI to become increasingly welcomed. In other words, its very success creates trust. As that trust is earned on relatively simple and safe repetitive tasks and seeing AI succeed in more complex jobs, it will be asked to take on ever more sophisticated tasks.
As trust builds up, the critical risk is that we become so dependent on AI that the authorities cannot exercise control without it. In a sense, AI optimises to become irreplaceable. By then, turning the AI engine off may be impossible because it performs vital functions, and the risk of disastrous outcomes might be deemed unacceptably high. Eventually, we risk becoming dependent on a system for critical analysis and decisions we don’t entirely, or even partially, understand.